Calculate the pH of a 1.41×10-2 M solution of the decongestant ephedrine hydrochloride if the pKb of ephedrine (its conjugate base) is 3.86.
Answer #1

ephedrine hydrochloride is salt of weak base and strong base

for such salts

pOH = 1/2 [pKw + pKb + log C]

pKb = 3.86

C = 0.0141 M

pOH = 1/2 [14 + 3.86 + log 0.0141]

pOH = 8.00

pH = 14 - 8.0

pH = 6.0